# LLM Query Scheduling with Prefix Reuse and Latency Constraints

### v1

@@ 4.2 TTFT Improvement from $k$-LPM
LPM: Label the queries so that $1,2,\dots,n$ are in ascending order of arrival times, i.e. $t_{1}\leq t_{2}\leq\cdots\leq t_{n}$. Let $\sigma(\cdot)$ be the permutation specifying the processing order under LPM. We show that with high probability, some query among the earliest arrivals (say indices $[q]$) is processed in a very late position ($>j$).
@@
$\bigl{\{}\exists\,i\in[q]:\sigma(i)>j\bigr{\}}\;=\;\bigl{\{}\text{some earliest-$q$ arrival is not processed among the first $j$ positions}\bigr{\}}.$
@@
$\begin{cases}(u+d)\,,&\text{if the user prefix differs from that of query $i-1$,}\\[6.0pt] d\,,&\text{if query $i$ has the same user prefix as query $i-1$.}\end{cases}$
@@
$I_{i}~{}\;=\;\begin{cases}1,&\text{if queries $i$ and $i-1$ share the same user prefix},\\ 0,&\text{otherwise}.\end{cases}$
Note that $\mathbb{E}[I_{i}]=\frac{k-1}{n-1}$ for all $i=2,...,n$, since the probability that the $i$ and $(i-1)$-th queries share the same (user) prefix is $\frac{k-1}{n-1}$. Then, the time needed to process the entire queue is given by:
@@
Corollary 3.
For any values of $s$, $k$, $u$, and $d$ such that $0<s<u$ and $k>1$, and for any $\delta\in(0,1)$, there exists $n_{0}\in\mathbb{N}$ such that $k$-LPM achieves a lower maximum TTFT that LPM and FCFS simultaneously on the regular arrival shuffled queue (Definition 4) for any value of $n\geq n_{0}$ with probability at least $1-\delta$.
This result holds for any value of $c_{\mathrm{attn}}\geq 0$ in the computational model of Definition 2.
@@
5 Experiments
5.1 Experimental Setup
In our experiments, we use the Llama-3.1-8B-Instruct model [4] with tensor parallelism across eight A100 GPUs. We run the experiment using the SGLang v0.4.1 serving framework. In particular, we evaluate the timing metrics using SGLang’s serving benchmark utility [13] and only modify the benchmarking dataset. We implement the $k$-LPM algorithm as an extension to the current LPM implementation in SGLang. Finally, we construct the dataset used for benchmarking by sampling four prompts with shared user history from the 8k context length prompts described in [17] for $2100$ prompts in total. We then randomly shuffle the ordering of these prompts and use $100$ for warm up of the benchmarking server.
5.2 Performance of $k$-LPM
We measure P99 TTFT vs. request rate for varying values of $k$ in the $k$-LPM algorithm. Our key observation (Figure 2) is that setting the hyperparameter to $k=2$ achieves reduced P99 TTFT across a wide range of request rates compared to FCFS and LPM. Note that SGLang’s benchmarking utility uses a Poisson arrival process, so “request rate” refers to the average number of requests per second.
@@ 5.2 Performance of $k$-LPM
Figure 2: We measure P99 TTFT versus request rate for five values of the hyperparameter $k$ on 2000 randomly shuffled prompts from the usecase described in [17]. Note that $k=1$ corresponds to FCFS and $k=\infty$ corresponds to LPM.
  • $k$-LPM performance remains robust under realistic models and prompt data, even when the theoretical assumptions are relaxed. Despite using a real LLM, an off-the-shelf serving framework, and a prompt dataset from an industrial use case, the observed scheduling behaviors closely match our theoretical predictions.
    First, in our experiments, prompt structures are only approximately tree-like. In practice, user histories exhibit negligible overlap, which does not noticeably affect $k$-LPM’s performance. Moreover, the near-random tie-breaking assumptions in Theorem 2 appear reasonable for these real-world prompts.
    Second, in the experiments, query processing begins as soon as the first query arrives rather than waiting for the entire query stream to finish. This implies that the assumption that processing begins at time $T\geq sn$ in Theorem 2 is inconsequential when the randomness in the order of user queries is on the order of the average queue size.
    Finally, relaxing the exact match between $k$ and the number of replicas still reduces TTFT. Although Theorem 2 assumes $k$ equals the number of user history replicas exactly, our experiments with four replicas of each user history show that smaller choices (e.g., $k=2$ or $k=3$) still yield substantial TTFT improvements.
6 Future Work
The primary objectives of this work were to formalize the LLM query scheduling problem in the context of RadixAttention and to develop a practical scheduling algorithm for latency-sensitive applications. This theoretical framework not only informs the design of new methods, but is particularly valuable given the current lack of empirical benchmarks for comparing scheduling strategies. However, further advancements in this area will necessitate empirical evaluation using real-world arrival patterns
The proposed problem formulation and theoretical results leave many interesting extensions open. One open question is whether there is an algorithm which returns a schedule satisfying a constraint on the $(1-p)$-th percentile TTFT in $\operatorname{poly}(1/p)$ time, i.e., a polynomial time approximation scheme. Additional directions of interest would be extending the computational model of Section 2 to handle distributional query streams or non-constant decoding length. Finally, generalizing the data generative model of Definition 4 may be interesting to capture other properties of real data.
Appendix A Percentile TTFT Constraint Approximation
Theorem 4 proves there is an algorithm that runs in polynomial time with respect to the query stream length that either returns a schedule satisfying constraint $T$ on the $(1-p)$-th percentile TTFT or certifies that no feasible schedules exists such that all queries satisfy the TTFT constraint $T$. Note that there is a necessary gap where the the positive case pertain the the percentile TTFT problem and the negative case certifies no schedule satisfies a constraint on the maximum TTFT. This gap is fundamental, since otherwise we could solve the problem for the maximum TTFT in by adding enough unsatisfiable queries and solving the percentile TTFT decision problem.
Theorem 4.
  1. 1.
  2. 2.
    Returns a schedule for $\mathcal{Q}$ such that the $(1-p)$-th percentile TTFT is at most $T$,
Proof.
At a high level we will prove that, if there exists a schedule for $\mathcal{Q}$ satisfying the TTFT constraint $T$ for all queries, then there exists a subset $\mathcal{Q}^{\prime}\subset\mathcal{Q}$ of size $(1-p)n$ that also satisfies the constraint $T$ for all queries. By the contrapositive of the statement, if $\mathcal{Q}^{\prime}$ does not exist, then no such schedule exists for $\mathcal{Q}$. We show that $\mathcal{Q}^{\prime}$ can be constructed by decomposing $\mathcal{Q}$ into $\Theta(n)$ subproblems, each of which can be solved in time independent from $n$, thereby providing an algorithm that is tractable with respect to $n$.
First, note that if there exists a feasible schedule for $\mathcal{Q}^{\prime}$ under a uniform TTFT constraint $T$, then there exists a schedule for $\mathcal{Q}$ where the $(1-p)$-th percentile TTFT is at most $T$. This schedule can be constructed by following the schedule for all queries in $\mathcal{Q}^{\prime}$ and then processing the remaining queries afterwards. By the contrapositive statement, if there does not exist such a schedule for $\mathcal{Q}^{\prime}$, then there does not exist such a schedule for $\mathcal{Q}$.
  1. 1.
    Let the possible cache initialization of $\mathcal{Q}_{k}^{\prime}$ be the feasible end queries of $\mathcal{Q}_{k-1}^{\prime}$ or the empty string if $k=1$. Then, in order of $k=1,...,d$, compute all possible pairs of a cache initialization and last query processed in $\mathcal{Q}_{k}$ where a feasible schedule satisfying the constraint $T$ exists.
  2. 2.
  3. 3.
  1. 1.
  2. 2.
  3. 3.

### v3

@@ Appendix A Percentile TTFT constraint approximation
At a high level we will prove that, if there exists a schedule for $\mathcal{Q}$ satisfying the TTFT constraint $T$ for all queries, then there exists a subset $\mathcal{Q}^{\prime}\subset\mathcal{Q}$ of size (1p)n(1-p)n that also satisfies the constraint $T$ for all queries. By the contrapositive of the statement, if $\mathcal{Q}^{\prime}$ does not exist, then no such schedule exists for $\mathcal{Q}$. We show that $\mathcal{Q}^{\prime}$ can be constructed by decomposing $\mathcal{Q}$ into $\Theta(n)$ subproblems, each of which can be solved in time independent from $n$, thereby providing an algorithm that is tractable with respect to $n$.
First, note that if there exists a feasible schedule for $\mathcal{Q}^{\prime}$ under a uniform TTFT constraint $T$, then there exists a schedule for $\mathcal{Q}$ where the (1p)(1-p)-th percentile TTFT is at most $T$. This schedule can be constructed by following the schedule for all queries in $\mathcal{Q}^{\prime}$ and then processing the remaining queries afterwards. By the contrapositive statement, if there does not exist such a schedule for $\mathcal{Q}^{\prime}$, then there does not exist such a schedule for $\mathcal{Q}$.
  1. 1.
    Let the possible cache initialization of $\mathcal{Q}_{k}^{\prime}$ be the feasible end queries of $\mathcal{Q}_{k-1}^{\prime}$ or the empty string if $k=1$. Then, in order of k=1,,dk=1,...,d, compute all possible pairs of a cache initialization and last query processed in $\mathcal{Q}_{k}$ where a feasible schedule satisfying the constraint $T$ exists.
  2. 2.
  3. 3.
  1. 1.
  2. 2.
  3. 3.
Appendix B Proofs for Section 4
B.1 Proof of Theorem 2
Proof.
LPM: Label the queries so that $1,2,\dots,n$ are in ascending order of arrival times, i.e. $t_{1}\leq t_{2}\leq\cdots\leq t_{n}$. Let $\sigma(\cdot)$ be the permutation specifying the processing order under LPM. We show that with high probability, some query among the earliest arrivals (say indices $[q]$) is processed in a very late position (>j>j).
@@ B.1 Proof of Theorem 2
[…]
@@
[…]
@@
[…]
Note that $\mathbb{E}[I_{i}]=\frac{k-1}{n-1}$ for all i=2,,ni=2,...,n, since the probability that the $i$ and (i1)(i-1)-th queries share the same (user) prefix is $\frac{k-1}{n-1}$. Then, the time needed to process the entire queue is given by:
@@
B.2 Proof of Corollary 3
@@
Appendix C Additional experiment results
C.1 Real prompt distribution latency metrics
In this section, we provide additional plots for latency metrics corresponding to the experiment of Figure 2.
Note that the relative performance of the scheduling algorithms is accurately predicted by Theorem 2 and surrounding discussion. That is, LPM (equivalent to $k$-LPM with $k=\infty$) attains the best median TTFT as it maximizes prefix reuse. However, $k$-LPM with intermediate values of $k$ attains nearly the same median TTFT and better P99 TTFT by ensuring that no queued query is waiting for too long.
(a) Median TTFT
(b) P90 TTFT
(c) P95 TTFT
(d) P99 TTFT
@@ C.1 Real prompt distribution latency metrics
Figure 3: We measure P50, P90, P95, and P99 TTFT versus request rate for five values of the hyperparameter $k$ on 2000 randomly shuffled prompts from the usecase described in 360brew. Note that $k=1$ corresponds to FCFS and $k=\infty$ corresponds to LPM.
C.2 Varying prefix length ratio
In this section, we provide experiments which vary $u$ and $d$ in the Regular Arrival Shuffled Queue of Definition 4555Note that we generate the prefix and suffix as random strings of the specified length, which results in non-zero negligible reuse between unique prefix or suffix substrings.. We set $n=2000$, $k=4$ (the number of replications per prefixes), and measure at 80 requests per second. For this experiment, we use a single H100 GPU and the Meta-Llama-3.2-1B-Instruct model dubey2024llama.
Table 3 shows latency results for $k$-LPM with $k$ equal to $1$, 44, and $\infty$ respectively, while varying the ratio of the replicated prefix to the constant prompt length (see Definition 4). These results match the relative performance of the settings of $k$ predicted by Theorem 2. That is, LPM consistently results in lower median TTFT due to greater prefix reuse. However, when the ratio of the prefix length to the prompt length ($u$ to $u+d$) is small, FCFS achieves lower TTFT at high percentiles by avoiding any query from waiting for too long. As the the ratio of the prefix length to the prompt length increases, the impact of better prefix reuse becomes greater, and so LPM performs better even regarding higher percentile TTFT. Finally, $k$-LPM with intermediate $k$ never performs simultaneously worse than FCFS and LPM, while it performs better than both in the intermediate regime.
Table 1: TTFT timings in seconds while varying the ratio of the repeated prefix to a constant total string length.
(a) FCFS
$u$ $d$ Median P90 P95
$1000$ $5000$ $3466.21$ $5604.33$ $5843.80$
$3000$ $3000$ $509.93$ $1040.45$ $1081.62$
$5000$ $1000$ $20.15$ $524.91$ $564.44$
(b) k-LPM
$u$ $d$ Median P90 P95
$1000$ $5000$ $1940.35$ $8610.27$ $9128.23$
$3000$ $3000$ $177.40$ $821.41$ $1647.49$
$5000$ $1000$ $19.61$ $273.55$ $447.37$
(c) LPM
$u$ $d$ Median P90 P95
$1000$ $5000$ $677.07$ 11 297.5311\,297.53 14 733.4514\,733.45
$3000$ $3000$ $176.32$ $1245.15$ $2821.72$
$5000$ $1000$ $20.70$ $216.88$ $565.92$